OpenStudy (anonymous):

how do you factor a(x+y)-(x+y)^2 (please explain how to do it. thankyou)

OpenStudy (anonymous):

(x+y)(a-(x+y))=(x+)(a-x-y)=(x+y)(-x-y+a)

OpenStudy (mathstudent55):

\(a(x+y)-(x+y)^2\) Think of x + y as being just z; let z = x + y, then you have \(az - z^2\) With the expression with z, it's easy to see that the common factor is z, so you factor out a factor of z: \(z(a - z)\) Now you can substitute back x + y for z, giving you: \((x + y)[a - (x + y)]\) Now you can simplify the second term, giving you: \((x + y)(a - x - y)\)

OpenStudy (anonymous):

Thanks :) It took me a month to go online again, but thanks :DD